OpenStudy (anonymous):

Let's see how you're at calculus ;). Differentiate the following function:

OpenStudy (anonymous):

\[f(x)=\ln(x^2+10)\]

OpenStudy (anonymous):

\[\frac{ d }{ dx }\ln x=\frac{ 1 }{ x}\]

OpenStudy (anonymous):

good job, surjthayer!

OpenStudy (anonymous):

Next medal goes for the person who can explain it.

OpenStudy (anonymous):

I wonder if this is your problem or you actually wondering if someone can do it.

OpenStudy (anonymous):

Not you Surj, you had your medal cheeky. ;)

OpenStudy (anonymous):

If people can actually do it.

OpenStudy (anonymous):

\[\huge f(x) = \ln(x^2+10)\] \[\huge f'(x) = \frac{ 1 }{ 10+x^2 }\times (10+x^2)'\] \[\huge = \frac{ 2x }{ x^2+10 }\]

OpenStudy (anonymous):

correct. took natural logarithms on bothside, differentiated with respect to x, and found the derivative. splendid job iambatman as well as surjthayer

OpenStudy (anonymous):

Yes (d/dx) [f(u(x)] = [1/u(x)](du/dx]

OpenStudy (anonymous):

Next challenge or no-go?

OpenStudy (anonymous):

Next challenge or no-go?

OpenStudy (anonymous):

Now opening new question.

OpenStudy (anonymous):

Now opening new question.